Write the equation of the line that has a y-intercept of (0,-6) and is perpendicular to the line 2x - 6y = 8.
Answer by stanbon(75887)

Write the equation of the line that has a y-intercept of (0,-6) and is perpendicular to the line 2x - 6y = 8.
-------------
The slope of the given line is 2/6 = 1/3
--
So the slope of the perpendicular line is -3
---
The y-intercept of the perp line is (0,-6)
-----------------
Equation is y = -3x - 6
===========================
Cheers,
Stan H.